40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 25 19:42:12 2014, ericq, Update, General, Vent progress 
    Reply  Fri Sep 26 12:25:34 2014, ericq, Update, General, Vent progress 
       Reply  Fri Sep 26 16:10:14 2014, ericq, Update, General, Vent update 
          Reply  Fri Sep 26 17:13:39 2014, ericq, Update, General, Vent update 
             Reply  Mon Sep 29 12:47:51 2014, ericq, Update, General, Vent update 
                Reply  Mon Sep 29 17:10:51 2014, ericq, Update, General, Vent update 
                   Reply  Mon Sep 29 18:12:24 2014, ericq, Update, General, Vent update 
                      Reply  Tue Sep 30 11:53:29 2014, ericq, Update, General, Vent update 
                         Reply  Tue Sep 30 16:18:57 2014, Steve, Update, VAC, pump down #78 has started pd78hastarted.png
                            Reply  Tue Sep 30 18:02:53 2014, Steve, Update, VAC, pump down #78 stops at 320 Torr 3h.png
                               Reply  Wed Oct 1 10:21:19 2014, Steve, Update, VAC, pump down #78 stops at 3.5 Torr pd78secondStop.png
                                  Reply  Wed Oct 1 16:26:53 2014, Steve, Update, VAC, pump down #78 completed pd78completedin25hrs.png
                                     Reply  Thu Oct 2 09:23:23 2014, Steve, Update, VAC, vent 77 vent77SRMmoved.png
                                        Reply  Fri Oct 3 10:10:37 2014, Steve, Update, VAC, cold cathode gauge reading switched CC1switched.png
                                           Reply  Fri May 8 14:30:09 2015, Steve, Update, VAC, CC1 cold cathode gauges are baked now cc1bake95C.jpg
                                              Reply  Tue May 12 14:57:52 2015, Steve, Update, VAC, CC1 cold cathode gauges are baked now 
Message ID: 10559     Entry time: Thu Oct 2 09:23:23 2014     In reply to: 10557     Reply to this: 10563
Author: Steve 
Type: Update 
Category: VAC 
Subject: vent 77 

Quote:

 

 Pump down reached "vacuum normal" state. IFO _P1 pressure 1e-4 torr

PSL shutter is opened.

 IFO_P1 pressure 1.6e-5 torr after 6 days at atm

 

PS: PSL sliding door 11 was left open overnight. The PSL particle count will reach room counts in 20 seconds at low speed of HEPA

Attachment 1: vent77SRMmoved.png  68 kB  | Show | Show all
ELOG V3.1.3-